Compare the electrostatic force between an electron and proton separated by 0.530x10^-10m with the gravitational force between them. This distance is their average separation in a hydrogen atom (r=0.530x10^-10m)
Answer #1

The electrstatiel force between two charges q1 and q2 separated by distance r is-

F​​​​​​e=kq1q2/r​​​​​2

where k=9x109N-m2/c​​​​​​2​​​

q1=charge on electron =-e

q2=charge on proton=+e

r=0.53x10-10m

q1 and q2 are opposite sign ,so force is attractive.

Where e=1.6x10-19c

F​​​​​​e=9x109x(1.6x10-19)2/(0.53x10-10)2=8.2x10-8Newton

mass of electron m1=9.11x10-31kg

Mass of proton m2=1.67x10-27kg

The gravitational force-

F​​​​​​g​​​​​=Gm1m2/r​​​​​​2

Where G is gratitational constant=6.674x10-11N-m2/kg​​​​​2

F​​​​​​g=​​​​(6.674x10-11x9.11x10-31x1.67x10-27)/(0.53x10-10)2

F​​​​​​g​​​​​​​​​​=3.6x10-47N

F​e/F​​​​​​g =(8.2x10-8)/(3.6x10-47)=2.28x1039N

so electrostatic force is 2.28x1039 times of gravitational force.